it has been found that 80%of the geometry students pass the chapter test on conguent triangles. If a class of 15 students, find the probability that a least one student will fail
1
Expert's answer
2013-03-19T09:09:31-0400
If there are 15 students in a class, then

15·0.8 = 12

students pass the chapter test on congruent triangles and the probability that a student will pass the chapter test is

p = 12/15 = 4/5 = 0.8.

Therefore, the probability that a least one student from 15 will fail is

P = 1 - (no one student will fail) =

= 1 - 0.8^15 ≈ 0.965.